OpenStudy (anonymous):

Solve the system. y = x + 2 y = -x + 0

OpenStudy (whpalmer4):

You can add the two equations together: y = x + 2 y =-x + 0 --------- 2y = 0x + 2 y = 1 Now use y = 1 in one of the original equations to find x. If you had two equations where none of the terms had the same coefficients, you can multiply both sides of an equation by the same number to make it happen. In this case we had x and -x, so we added, but if both had had just x, we could have subtracted instead of adding, or multiplied one by -1 on both sides.